OpenStudy (anonymous):

can someone help me please x^1/4 =16

OpenStudy (anonymous):

Raise each side to the 4th power. (x^(1/4))^4 = x So, just raise6 to the 4th power.

OpenStudy (anonymous):

16 to the 4th power, sorry.

OpenStudy (anonymous):

x^1/4 = 16 \[\sqrt[4]{x} = 16\] \[x = 16^4\]

OpenStudy (anonymous):

So, what is (16)(16)(16)(16) ?

OpenStudy (anonymous):

are you serious? It was that easy?

OpenStudy (anonymous):

Sometimes we make problems harder than necessary. We all do that at times.

OpenStudy (anonymous):

thank you both so much

OpenStudy (anonymous):

uw!